OpenStudy (anonymous):

A zero-order reaction has a constant rate of 4.80×10−4M/s . If after 60.0 seconds the concentration has dropped to 7.00×10−2M , what was the initial concentration?

OpenStudy (anonymous):

help!

OpenStudy (anonymous):

Concentration at time t = -(Constant rate x time t) + Initial concentration

OpenStudy (anonymous):

7.00×10^(-2)=-(4.80×10−4)(60) + Initial concentration using algebra, we get (7.00×10^(-2)) + ((4.80×10−4)(60)) = Initial concentration = 9.88^(-2) M :))